OpenStudy (anonymous):

okay really confused now!!! 31^2=square root of 177+b^2????? huh???!!!

OpenStudy (anonymous):

Is this the original problem?

OpenStudy (anonymous):

yes I typed it right from the book were working

OpenStudy (anonymous):

Ok

OpenStudy (anonymous):

So what is 31^2?

OpenStudy (anonymous):

31^2=961 961-177 = 784 sqrt(784)=28 b=28

OpenStudy (anonymous):

Then once you figure that out, set both sides to the power of 2 , to get rid of the sqrt sign

OpenStudy (anonymous):

Lol, ^ what he said then.

OpenStudy (anonymous):

lol ok

OpenStudy (anonymous):

Wait hold on.. is the 177+b^2 all inside the parenthesis? or seperate?

OpenStudy (anonymous):

ok, sunshine, do you know what Pythagorean theorem is used for?

OpenStudy (anonymous):

Is the problem like this \[31^2 = \sqrt(177)+b^2 or 31^2=\sqrt(177+b^2)\]

OpenStudy (anonymous):

but don't you square root the 177 first???

OpenStudy (anonymous):

Depends. That is why you need to tell me which one the book is saying. Is 177 the only number inside the square root?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

Okay based on what you said the problem is it is similar to what bpw12 did: 31^2=961 961-sqrt177 = 947.696 b= sqrt(947.696)=30.785 b=28

OpenStudy (anonymous):

ok ill retype the problem exactly [31^2=sqrt{177}+b^2

OpenStudy (anonymous):

Yeah, then it is what I said to do then.

OpenStudy (anonymous):

ok do i square root the 177 at the end???

OpenStudy (anonymous):

Nope. After you subtract sqrt {177} from both sides you have: 31^2-sqrt{177}=b^2 947.696=b^2 sqrt(947.696)=sqrt(b^2) Now you may or may not learned this next part, but technically there will be two solutions for b now that I think about (sorry mind is being slow tonight). So b=30.785 or b=-30.785 P.S. Ignore the b=28 from my post from like 8 minutes ago

OpenStudy (anonymous):

ok I'm skipping this prob ill ask my teacher tomorrow